OpenStudy (xmoses1):

The slope of the line passing through the points (-1, 5) and (3, 5) is no slope 0 5/4

OpenStudy (xmoses1):

I chose No Slope

jimthompson5910 (jim_thompson5910):

use the slope formula m = (y2-y1)/(x2-x1)

OpenStudy (xmoses1):

ok one sec

jimthompson5910 (jim_thompson5910):

In this case, (x1, y1) = (-1, 5), which means x1=-1 and y1= 5 and (x2, y2) = (3, 5), which means x2=3 and y2= 5 ------------------------------------------------------- So x1=-1 y1= 5 x2=3 y2= 5

OpenStudy (xmoses1):

I got m=0/4...

jimthompson5910 (jim_thompson5910):

reduce that to get ???

OpenStudy (xmoses1):

4 is the slope?

jimthompson5910 (jim_thompson5910):

no

jimthompson5910 (jim_thompson5910):

0/4 is the slope but you have to reduce

OpenStudy (xmoses1):

...

OpenStudy (xmoses1):

0/4 doesn't = 4?

OpenStudy (xmoses1):

Oh wait 4/0 would = 4 right?

OpenStudy (xmoses1):

so it has a slope of 0

jimthompson5910 (jim_thompson5910):

4/0 is undefined since you CANNOT divide by zero

jimthompson5910 (jim_thompson5910):

0/4 is 0 since 0 divided by any number (but 0) is 0 in other words 0/x = 0 where x is not 0

OpenStudy (xmoses1):

ok so the slope is 0

jimthompson5910 (jim_thompson5910):

correct

OpenStudy (xmoses1):

ok thank you for explaining too

jimthompson5910 (jim_thompson5910):

yw

OpenStudy (xmoses1):

oh could you just tell me if i have the right answer for this one question real quick? The slope of the line whose equation is 3y = 2x - 3 is 2 2/3 3/2 My answer was 2, am i right?

jimthompson5910 (jim_thompson5910):

that's incorrect unfortunately

jimthompson5910 (jim_thompson5910):

you have to get the equation into y = mx+b form m will be the slope

OpenStudy (xmoses1):

huh that might explain why i got a 67 on this quiz lol

jimthompson5910 (jim_thompson5910):

that's ok, you can improve and learn from your mistakes

OpenStudy (xmoses1):

yeah its a good thing that i have OS because my home school is not very good at teaching

OpenStudy (xmoses1):

so y=3/2 x-3

jimthompson5910 (jim_thompson5910):

not quite

jimthompson5910 (jim_thompson5910):

you have to divide both sides by 3

OpenStudy (xmoses1):

oh so y= 2/3x-1

jimthompson5910 (jim_thompson5910):

yes, so the slope is _______

OpenStudy (xmoses1):

2/3 lol

jimthompson5910 (jim_thompson5910):

correct

OpenStudy (xmoses1):

thank you!

jimthompson5910 (jim_thompson5910):

you're welcome